According to Nielsen Media Research, the (population) average number of hours of TV viewing per household per week in the United States is 50.4 hours. Suppose the (population) standard deviation is known to be 11.8 hours. A random sample of 42 U.S. households is taken. What is the probability that the sample mean is more than 52 hours?



1
Expert's answer
2021-05-07T09:16:01-0400

Let X denotes number of hours of TV viewing per household.

X has mean = 50.4

We have sample standard deviation = 11.8

n = 42

Probability that the sample mean is more than 52 hours

"= P(\bar X>52)"

"= P(Z>\dfrac{52-50.4}{\dfrac{11.8}{\sqrt 42}})"


"= P(Z> 0.87)"

"= 0.5 - P(0<Z<0.87)"

"= 0.5-0.3078"

"= 0.19"